# Concatenate, Fine-tuning, Re-training: A SAM-enabled Framework for Semi-supervised 3D Medical Image Segmentation

### v1

@@ V Fine-tuning Module
Our fine-tuning module performs fine-tuning on the vision foundation model. As one of the most popular universal image segmentation models, SAM serves as the default setting of fine-tuning module in our CFR, denoted as . We strip away all the task prompts, and perform automatic segmentation during inference. Our framework is not limited to a specific fine-tuning strategy, which could be used in different strategies. We uniformly denote the fine-tuning loss :
@@
We notice that concatenating 2D slices may result in organs appearing in fixed regions, guiding SAM to capture these similarities. As shown in Figure 6, taking the spleen as an example, since the upsampling fine-tuning method predicts slice by slice, it confuses the spleen (red) into the left kidney (blue) between adjacent slices, even when the target is located in the same region across different slices. In contrast, our concatenate strategy preserves the spatial relationship between slices and could leverage the information of the same target from neighboring slices. To verify the effectiveness of pseudo-labels on different datasets, we visualize the fine-tuning module prediction in Figure 7. The mask decoder outputs a large-sized 2D mask, which is subsequently restored to a 3D volume as a pseudo-label for the SSL module.
VI Re-training SSL Module
@@
Self-training: Self-training [39] involves three iteratively steps: Firstly, a teacher network is initially trained on . Secondly, it makes predictions on to obtain . Thirdly, a student network retrained on the union set . Among them, steps 2 and 3 are iteratively performed in alternation. The unsupervised loss can be formulated as:
where and denote the weights of student and teacher networks, and means parameters are fixed.
@@
Mean Teacher: Mean teacher (MT) [13] consists of a student network and a teacher network, and teacher network weights are updated with the exponential moving average (EMA) of student network weights. The unsupervised loss is the consistency regularization between the two networks:
where can be a labeled sample or an unlabeled sample . and denote the input data with different perturbations of student and teacher networks respectively.
@@
ACMT: ACMT [9] improves MT to the ambiguity-consensus mean teacher model, encouraging consistency between the student’s and the teacher’s predictions at the identified ambiguous regions. The unsupervised consistency loss is redesigned as ambiguity-masked error:
where is the binary ambiguity indicator, and and are the predictions of two networks at the -th voxel.
@@
MagicNet: MagicNet [8] introduces a data augmentation strategy based on MT. First, a pair of labeled and unlabeled samples are mixed into two shuffled cubes. Next, small cubes and mixed cubes are fed into the segmentation network, and finally recovery the mixed cubes. The unsupervised loss function consists of two parts:
where is the small cubes location reasoning loss, and is the mix-recovered image consistency loss defined in [8].
@@
The supervised loss consists of the joint cross-entropy loss and Dice loss for Self-training, Mean Teacher, and ACMT methods, and for MagicNet, the supervised loss includes labeled image segmentation loss, location reasoning loss, and mix-recovered loss.
VII Summary
The training procedure of our CFR framework is summarized in Algorithm 1. During fine-tuning, the 3D labeled volumes are first transformed into larger-sized 2D images by the concatenation module. Then train the fine-tuning module which is updated by minimizing the supervised fine-tuning loss. During re-training, the fine-tuning module annotates unlabeled samples and provides pseudo-labels to re-training module. The SSL re-training module is trained by incorporating the supervised information from labeled images and the pseudo-label consistency from unlabeled images.
Our CFR framework maintains the same parameter level as the mainstream segmenter and ensures computational efficiency. The fine-tuning module CFR with the LoRA strategy has the same parameter size as the foundation model (SAM). The re-training module has a parameter scale to the mainstream segmenters like V-Net [26]. During inference, we discard fine-tuning module and retain only re-training module. The results of new samples are directly predicted using .
VIII Experiments
The experiments are conducted on two single target datasets (i.e., LA and BraTS) and two multiple target datasets (i.e., BTCV and MACT). We perform experiments of semi-supervised segmentation with moderate annotations and scarce annotations on each dataset. Subsequently, we further analyze the pseudo-labels generated by the fine-tuning module, as well as the effectiveness and compatibility of the retraining module, and conduct experiments on different labeled samples.
VIII-A Datasets
BrsTS Dataset. The dataset contains preoperative MRI (with modalities of T1, T1Gd, T2 and T2-FLAIR) of 335 glioma patients from the BraTS 2019 challenge [48, 49, 50], where 259 patients with high-grade glioma and 76 patients with low-grade glioma. Following [9], we only use T2-FLAIR images with the same data split and pre-processing procedures for fair comparison.
@@ VIII-A Datasets
VIII-B Experimental Settings
In this paper, all the experiments are implemented in Pytorch on the NVIDIA GeForce RTX 3090/4090TI GPU. For foundation model SAM, we conduct all the experiments based on the “ViT-B” version. We adopt LoRA finetuning and the rank of LoRA is set to 4 for efficiency and performance optimization. In our experiments, we follow the current popular setting in designing mask decoder, i.e., SAMed [6] with LoRA, which modifies the segmentation head to generate masks of each class in a deterministic manner and aggregate them to final segmentation map.
For the semi-supervised network, it is trained by the SGD optimizer with an initial learning rate of 0.01. For LA and BraTS datasets, we follow the training strategy of [12, 9]. And we employ four measurements to quantitatively evaluate the segmentation performance, including Dice, Jaccard, the average surface distance (ASD), and the 95% Hausdorff Distance (HD). For BTCV and MACT datasets, we follow the implementation details of [8] and Dice as an evaluation metric for multi-organ segmentation. For fair comparison, we use official reported of various baselines [12, 9, 8].
VIII-C Results with Moderate Annotations
VIII-C1 Single Target Segmentation
We compare with current SOTA methods on LA dataset and BraTS dataset, including V-Net [26], MT [13], UA-MT [12], CCT [14], CPS [15], DTC [43], ICT [44], CPCL [45], URPC [46], BCP [10], MCCauSSL [17] and ACMT [9], and the results are shown in Table I and Table II. We select the basic MT [13] network and method ACMT [9] with the current highest performance, and respectively provide pseudo-labels of adapted SAM to assist training. Our CFR framework with MT and ACMT as the retraining module is denoted as CFR and CFR, respectively. The results show that CFR framework brings substantial improvements in both MT and ACMT methods, approaching the performance of fully-supervised segmentation.
VIII-C2 Multiple Target Segmentation
Table V and Table VI show the results on BTCV and MACV datasets, respectively. We compare our method with V-Net [26], MT [13], UA-MT [12], CPS [15], ICT [44], SS-Net [53], SLC-Net [54], and MagicNet [8]. We employ CFR framework to guide the classical MT [13] and the SOTA MagicNet [8] methods, i.e., CFRMT and CFRMagicNet. By leveraging the CFR framework, both methods achieve significant performance improvements. On the BTCV dataset, CFRMagicNet even surpasses the results obtained with fully-supervised learning.
@@ VIII-C2 Multiple Target Segmentation
| CFR (Ours) | 49.08 (6.18) | 89.57 | 74.00 | 33.21 | 16.55 | 91.05 | 46.91 | 25.84 | 15.47 |
VIII-D Results with Scarce Annotations
VIII-D1 Single Target Segmentation

### v2

@@ III-B2 Fine-tuning Module
Our fine-tuning module performs fine-tuning on the vision foundation model. As one of the most popular universal image segmentation models, SAM serves as the default setting of our fine-tuning module, denoted as . We strip away all the prompts and perform automatic segmentation during inference. Our framework is not limited to a specific fine-tuning strategy, which could be used in different strategies. We uniformly denote the fine-tuning loss :
@@
LoRA offers a practical and effective approach for fine-tuning across domains, and the combination of LoRA with our specific stitching strategies tailored for medical images ensures that our model could handle the substantial differences between natural and medical image data well.
We notice that stitching 2D slices may lead to organs appearing in surrounding and similar regions, guiding SAM to capture these similarities. Our stitching strategy preserves the spatial relationship between slices and could leverage the information of the same target from neighboring slices effectively. For example, as shown in Fig. 6, the same semantic class (e.g., spleen) appears in a similar area (upper left) across three adjacent slices. The upsampling fine-tuning method, which predicts slice-by-slice, confuses the entire spleen (red) into the left kidney (blue) in the middle slice. In contrast, our stitching fine-tuning method could correctly identify the spleen in all three slices. By incorporating inter-slice information, our method effectively guides SAM in recognizing the same organ on different slices. In contrast, our stitching strategy preserves the spatial relationship between slices and could leverage the information of the same target from neighboring slices.
To verify the effectiveness of pseudo-labels on different datasets, we visualize the fine-tuning module prediction of single and multiple target datasets, taking LA [15] and BTCV [44] datasets as examples, in Fig. 7. The mask decoder outputs a large-sized 2D mask, which is subsequently restored to a 3D volume as a pseudo-label for the SSL module.
III-B3 Re-training SSL Module
@@
Self-training: Self-training [43] involves three iteratively steps: Firstly, a teacher network is initially trained on . Secondly, it makes predictions on to obtain . Thirdly, a student network retrained on the union set . Among them, steps 2 and 3 are iteratively performed in alternation.
@@
Mean Teacher: Mean teacher (MT) [21] consists of a student network and a teacher network, and teacher network weights are updated with the exponential moving average (EMA) of student network weights. The unsupervised loss is the consistency regularization between the two networks.
@@
ACMT: ACMT [17] improves MT to the ambiguity-consensus mean teacher model, encouraging consistency between the student’s and the teacher’s predictions at the identified ambiguous regions.
@@
MagicNet: MagicNet [16] introduces a data augmentation strategy based on MT. First, a pair of labeled and unlabeled samples are mixed into two shuffled cubes. Next, small cubes and mixed cubes are fed into the segmentation network, and finally recovery the mixed cubes.
III-C SFR+ Framework
To further enhance our framework, we developed the extension version of SFR, named SFR+, as illustrated in Fig. 8. In the SFR framework, the fine-tuning module trained with labeled samples provides pseudo-labels of unlabeled samples uni-directionally to the re-training module. To more effectively extract and leverage information of unlabeled data from the two modules, our SFR+ introduces a confidence estimation strategy to distinguish between confident and uncertain samples. This approach enables selective optimization of both the fine-tuning and re-training modules.
III-C1 Confidence Estimation
SFR+ introduces confidence estimation to determine how to handle each unlabeled sample. We calculate the voxel-level average confidence for each unlabeled sample and classify them based on a threshold. For each unlabeled sample , the voxel-level average confidence is defined as:
where is the model prediction for voxel and indexes over the classes. is the number of voxels in . Samples are classified based on a threshold :
High-confident samples: If , the sample is considered confident and to update the fine-tuning module.
Low-confident samples: If , the sample is classified as uncertain and sent to the re-training module.
III-C2 Selective Training Strategy
In SFR+, selective learning in the fine-tuning and re-training modules allows for more effective handling of unlabeled samples. On one hand, high-confident samples are employed to update the fine-tuning module, ensuring that only reliable information from unlabeled data contributes to further refinement. On the other hand, low-confidence samples benefit from the pseudo-labels, which enable further improvements of the re-training module.
By selectively alternating updates between the two modules, SFR+ mitigates the risk of error propagation from inaccurate predictions that may arise due to the uni-directional transfer of pseudo-labels.
III-D Summary
The training procedure of our SFR framework is summarized in Fig. 2. During fine-tuning, the 3D labeled volumes are first transformed into larger-sized 2D images by the stitching module. Then fine-tuning module is updated by minimizing the supervised fine-tuning loss. During re-training, the fine-tuning module annotates unlabeled samples and provides pseudo-labels to re-training module. The SSL re-training module is trained by incorporating the supervised information from labeled images and the pseudo-label consistency from unlabeled images.
Our SFR and SFR+ framework ensures computational efficiency. The fine-tuning module SFR with the LoRA strategy has the same parameter size as the foundation model (SAM). The re-training module has a parameter scale to the mainstream segmenters like V-Net [30]. During inference, we discard fine-tuning module and retain only re-training module. The results of new samples are directly predicted using .
IV Experiments
The experiments are conducted on two single target datasets (i.e., LA [15] and BraTS [47]) and three multiple target datasets (i.e., BTCV [44], MACT [48], and AbdomenCT-1K [49]). We perform experiments of semi-supervised segmentation with moderate annotations and scarce annotations on each dataset. Subsequently, we further analyze the pseudo-labels generated by the fine-tuning module, as well as the effectiveness and compatibility of the retraining module, and conduct experiments on different labeled samples.
IV-A Datasets
BraTS Dataset. The dataset contains preoperative MRI (with T1, T1Gd, T2 and T2-FLAIR modalities) of 335 glioma patients from the BraTS 2019 challenge [47, 51, 52], where 259 patients with high-grade glioma and 76 with low-grade glioma. Following [17], we only use T2-FLAIR images with the same data split and pre-processing procedures for fair comparison.
AbdomenCT-1K Dataset. The AbdomenCT-1K dataset [49] is a diverse abdominal CT organ segmentation dataset, with more than 1,000 CT scans from 12 medical centers, including multi-phase, multi-vendor, and multi-disease cases. We follow the same pre-processing procedure as nnU-Net [53]. Similar to BTCV and MACT datasets, we divide the image into patches as input volumes.
IV-B Experimental Settings
In this paper, all the experiments are implemented in Pytorch on the NVIDIA GeForce RTX 3090/4090TI GPU. For foundation model SAM, we conduct all the experiments based on the “ViT-B” version. We adopt LoRA finetuning and the rank of LoRA is set to 4 for efficiency and performance optimization. In our experiments, we follow the current popular setting [10] in designing mask decoder, which modifies the segmentation head to generate masks of each class in a deterministic manner and aggregate to final segmentation map.
For the semi-supervised network, it is trained by the SGD optimizer with an initial learning rate of 0.01. For LA and BraTS datasets, we follow the training strategy of [20, 17]. And we employ four measurements to quantitatively evaluate the segmentation performance, including Dice, Jaccard, the average surface distance (ASD), and the 95% Hausdorff Distance (HD). For BTCV, MACT, and AbdomenCT-1K datasets, we follow the implementation details of Magicnet [16] and Dice as an evaluation metric for multi-organ segmentation. For fair comparison, we use official reported of various baselines [20, 17, 16]. In SFR+, we set the as 0.985.
IV-C Stitching Strategy Analysis
There is a performance trade-off between the resolution of individual slices and the number of slices stitched together into a single 1024 1024 image. Increasing the number of slices provides more context and spatial continuity. However, it also requires resizing the slices to fit into the fixed resolution, potentially resulting in a loss of fine-grained details. Therefore, how to consider both the slice resolution and the stitching scale becomes a critical issue.
@@ IV-C Stitching Strategy Analysis
In Fig. 9, we explored different stitching scales ( grid) to analyze this trade-off. On the LA dataset [15], when the stitch scale is set to , the large-sized 2D image represents the complete 3D volume, with each slice maintaining its original resolution without any downsampling. The results show that when the stitch scale is small, performance tends to be inferior. With a small scale, it is possible that the representation may not fully capture the volume context, limiting the model’s ability to leverage inter-slice information. By increasing the number of slices, the model benefits from more context, resulting in improved segmentation performance, particularly when the stitched image encompasses the entire 3D volume. When the resolution of individual slices is maintained, the large-sized 2D image represents the 3D volume without any downsampling, and the best results are achieved. We supplement this figure by further increasing the stitching scale, that is, reducing the resolution of each slice. However, it becomes counterproductive when further increasing the stitching scale—excessive downsampling of individual slices begins to degrade performance as the resolution becomes too low to capture critical anatomical details. In our method, each slice retains its resolution throughout the process, ensuring that no upsampling or downsampling occurs, which balances the need for spatial context with the preservation of resolution. The stitching process does not affect the resolution of the individual slices; rather, it combines them into a large-sized 2D image. As a result, the final 2D segmentation predictions are also aligned with the original slice resolution, preserving the fine-grained details essential for high-quality segmentation.
@@
| L / U | Method | AVG | Liver | Kidney | Spleen | Pancreas |
| 180 / 0 |
V-Net
[30] (LB) |
67.17 | 96.03 | 93.64 | 78.99 | 0* |
| 180 / 720 | MT [NIPS’17] [21] | 75.16 | 93.78 | 91.54 | 76.55 | 38.79 |
| SFR (Ours) | 88.73 | 96.10 | 94.55 | 93.63 | 70.65 | |
| SFR (Ours) | 90.00 | 96.11 | 94.50 | 95.19 | 74.19 | |
| MagicNet [CVPR’23] [16] | 90.54 | 96.38 | 94.54 | 95.17 | 76.08 | |
| SFR (Ours) | 91.41 | 96.70 | 95.01 | 96.02 | 77.90 | |
| SFR (Ours) | 91.70 | 96.66 | 95.20 | 96.06 | 78.88 | |
| 900 / 0 |
V-Net
[30] (UB) |
88.26 | 95.31 | 94.49 | 92.30 | 70.94 |
@@
In this work, our stitching module reorganizes a volume (3D raw image or 3D patch) into a image, enabling a complete representation of the volume within a single image. For the size of slices, different slices might vary a lot in different datasets. For large-scale medical datasets, such as multi-organ abdominal scans, the full 3D volume often has a very large voxel size (e.g., ). In 3D medical image segmentation approaches like 3D-UNet [57], V-Net [30], and nnU-Net [53], it is common practice to divide the volume into smaller 3D patches (e.g., ) for training and inference. In our approach, we follow a similar strategy during both the fine-tuning and retraining phases, using patches of consistent size. Before fine-tuning SAM, we stitch the 3D patches into a image, ensuring that the patch structure is preserved throughout the process.
Specifically, we divide the full 3D volume into consistent 3D patches, each maintaining an original resolution to avoid any loss of fine anatomical details. Each patch is then systematically stitched into a image to capture full spatial context. For the LA segmentation [15], we adopted the same strategy as works [20, 50, 17], where the patch size is set to where and . On the BraTS dataset [47], we followed the approach used in ACMT [17], using patch dimensions of . For the BTCV [44], MACT [48] and AbdomenCT-1K [49] datasets, following MagicNet [16], the patch size is set to . After fine-tuning, the predictions are restored to their 3D form through an inverse stitching transform, ensuring that the resolution and spatial integrity of the original 3D image are preserved. By carefully adjusting the patch size and stitching scale according to each dataset’s needs, our approach ensures robust performance across both single-target and multi-target segmentation tasks.
| L / U | Method | Dice | Jaccard | ASD | 95HD |
| 16 / 64 | MT [NIPS’17] [21] | 87.47 | 78.03 | 3.92 | 14.13 |
| SFR (Ours) | 89.09 (1.62) | 80.73 | 2.58 | 10.33 | |
| ACMT [MedIA’23] [17] | 88.50 | 79.56 | 3.58 | 14.71 | |
| SFR (Ours) | 89.49 (0.99) | 81.09 | 1.94 | 7.88 | |
| 1 / 79 | MT [NIPS’17] [21] | 44.89 | 29.85 | 18.86 | 44.78 |
| SFR (Ours) | 65.65 (20.76) | 53.64 | 9.82 | 29.43 | |
| ACMT [MedIA’23] [17] | 58.40 | 42.61 | 20.55 | 50.54 | |
| SFR (Ours) | 68.24 (9.84) | 57.40 | 12.38 | 30.76 |
In addition, we investigate a simplified strategy to ensure uniformity across different datasets and further validate the robustness of our framework. In this alternative approach, we processed all 3D medical images, regardless of their original size, by first cropping into patches. These patches are then directly stitched together to form a 2D image. Typically, the resolution of slices (both in height and width) in 3D medical images exceeds 96, though the number of slices (depth) may vary. For datasets where the number of slices is fewer than 96, we employ interpolation to achieve the desired depth. This approach provides a consistent standard for stitching, ensuring that the slices are uniformly processed before being fed into the model, without manual intervention. We have conducted experiments using this approach on the LA dataset with both moderate and scarce annotations, and the results are shown in Table VI. Notably, our method achieves over 89% Dice score on moderate annotations and outperforms SSL methods by a large margin on scarce annotations. This demonstrates that even with this simplified and fixed patch-size mechanism, the performance remains effective.
| L / U | Method | Dice | Jaccard | ASD | 95HD |
| 1 / 0 | V-Net
[30] (LB) |
17.99 | 12.93 | 19.66 | 44.58 |
| 1 / 79 | MT [NIPS’17] [21] | 29.68 | 18.19 | 18.63 | 42.58 |
| SFR (Ours) | 74.40 (44.72) | 61.47 | 6.25 | 25.90 | |
| SFR (Ours) | 81.03 (51.35) | 69.44 | 5.12 | 19.77 | |
| ACMT [MedIA’23] [17] | 72.64 | 58.33 | 10.42 | 33.22 | |
| SFR (Ours) | 76.76 (4.12) | 64.65 | 8.54 | 28.07 | |
| SFR (Ours) | 83.72 (11.08) | 72.80 | 4.54 | 18.60 |
| L / U | Method | Dice | Jaccard | ASD | 95HD |
| 1 / 0 | 3D U-Net
[57] (LB) |
73.74 | 61.44 | 13.81 | 37.07 |
| 1 / 249 | MT [NIPS’17] [21] | 63.52 | 50.59 | 20.59 | 47.54 |
| SFR (Ours) | 78.58 (15.06) | 66.56 | 7.16 | 23.43 | |
| SFR (Ours) | 79.04 (15.52) | 67.55 | 6.06 | 21.11 | |
| ACMT [MedIA’23] [17] | 63.32 | 51.41 | 9.28 | 31.71 | |
| SFR (Ours) | 78.47 (15.15) | 66.33 | 8.17 | 26.64 | |
| SFR (Ours) | 79.24 (15.92) | 67.56 | 6.72 | 22.88 |
| L / U | Method | AVG | Liver | Kidney | Spleen | Pancreas |
| 1 / 0 |
V-Net
[30] (LB) |
22.72 | 56.04 | 33.20 | 0* | 0.02 |
| 1 / 899 | MT [NIPS’17] [21] | 33.34 | 70.62 | 45.09 | 0.90 | 16.75 |
| SFR (Ours) | 57.89 | 72.09 | 66.06 | 71.33 | 22.10 | |
| SFR (Ours) | 65.12 | 77.40 | 71.35 | 77.27 | 34.45 | |
| MagicNet [CVPR’23] [16] | 54.19 | 75.26 | 62.44 | 48.82 | 30.23 | |
| SFR (Ours) | 64.32 | 75.04 | 68.08 | 76.12 | 38.03 | |
| SFR (Ours) | 66.16 | 79.91 | 71.63 | 70.91 | 42.18 |
IV-D Segmentation Results with Moderate Annotations
IV-D1 Single Target Segmentation
We compare with current SOTA methods on LA dataset and BraTS dataset, including V-Net [30], MT [21], UA-MT [20], CCT [22], CPS [23], DTC [50], ICT [54], CPCL [55], URPC [56], BCP [18], MCCauSSL [25] and ACMT [17], and the results are shown in Table I and Table II. We follow the same data partitioning and training strategies as in previous works [20, 18, 17, 16], ensuring that the comparison results are fair. In all experiments, we apply data augmentation techniques such as random cropping and flipping. It could be seen that BCP [18], CauSSL [25], and ACMT [17] perform well, which may be attributed to these methods are often well-designed for small datasets, where their sophisticated architecture may yield promising results. We select the basic MT [21] network and method ACMT [17] with the current highest performance, and respectively provide pseudo-labels of adapted SAM to assist training. Our SFR framework with MT and ACMT as the retraining module is denoted as SFR and SFR, respectively. The results show that our SFR framework brings substantial improvements in both MT and ACMT methods, approaching the performance of fully-supervised segmentation. Additionally, it is worth noting that SFR shows improvements in the Average Surface Distance (ASD) metric. In both datasets, our semi-supervised method SFR not only reduces the ASD but even surpasses the fully-supervised results. This improvement in ASD indicates a meaningful boost in boundary precision, which is critical for accurate medical image segmentation. Our SFR+ framework achieves a further improvement over SFR on both datasets. In Table I, it is clear that SFR+ improves the Dice score of ACMT from 90.31% to 91.00%, which is only 0.14% lower than the fully supervised result of 91.14%.
@@ IV-D1 Single Target Segmentation
| SFR (Ours) | 49.08 (6.18) | 89.57 | 74.00 | 33.21 | 16.55 | 91.05 | 46.91 | 25.84 | 15.47 | |
| SFR (Ours) | 54.05 (11.15) | 84.74 | 85.03 | 36.13 | 25.01 | 90.34 | 52.22 | 41.92 | 17.03 |
IV-D2 Multiple Target Segmentation
Table III, Table IV, and Table V show the results on AbdomenCT-1K, BTCV, and MACV datasets, respectively. We compare our method with V-Net [30], MT [21], UA-MT [20], CPS [23], ICT [54], SS-Net [58], SLC-Net [59, 60], and MagicNet [16]. We employ SFR framework to guide the classical MT [21] and the SOTA MagicNet [16] methods, i.e., SFR and SFR. By leveraging the SFR framework, both methods achieve significant performance improvements. On the BTCV dataset, SFR even surpasses the results obtained with fully-supervised learning. For the AbdomenCT-1K dataset, it could be observed that SFR improves performance across all four organs, boosting the average Dice score by 13.57% (88.73% vs. 75.16%) with 20% labeled data. In addition, our SFR+ framework demonstrates improvements across the three different datasets for moderate annotation scenarios, highlighting its robustness and effectiveness in segmentation.
The single target segmentation examples on LA dataset with 16 labeled data and multiple target segmentation examples on MACT dataset with 14 labeled data are shown in Fig. 11. Our SFR and SFR predictions align more accurately with ground truth masks, further validating the effectiveness. Our framework demonstrates its ability to adapt different data.
IV-E Segmentation Results with Scarce Annotations
IV-E1 Single Target Segmentation